### v2

# Improving Efficiency of Diffusion Models via Multi-Stage Framework and Tailored Multi-Decoder Architectures

Abstract
Diffusion models, emerging as powerful deep generative tools, excel in various applications. They operate through a two-steps process: introducing noise into training samples and then employing a model to convert random noise into new samples (e.g., images). However, their remarkable generative performance is hindered by slow training and sampling. This is due to the necessity of tracking extensive forward and reverse diffusion trajectories, and employing a large model with numerous parameters across multiple timesteps (i.e., noise levels). To tackle these challenges, we present a multi-stage framework inspired by our empirical findings. These observations indicate the advantages of employing distinct parameters tailored to each timestep while retaining universal parameters shared across all time steps. Our approach involves segmenting the time interval into multiple stages where we employ custom multi-decoder U-net architecture that blends time-dependent models with a universally shared encoder. Our framework enables the efficient distribution of computational resources and mitigates inter-stage interference, which substantially improves training efficiency. Extensive numerical experiments affirm the effectiveness of our framework, showcasing significant training and sampling efficiency enhancements on three state-of-the-art diffusion models, including large-scale latent diffusion models. Furthermore, our ablation studies illustrate the impact of two important components in our framework: (i) a novel timestep clustering algorithm for stage division, and (ii) an innovative multi-decoder U-net architecture, seamlessly integrating universal and customized hyperparameters.
1 Introduction
Figure 1: Overview of three diffusion model architectures: (a) unified, (b) separate, and (c) our proposed multistage architectures. Compared with (a) and (b), our approach improves sampling quality, and significantly enhances training efficiency, as indicated by the FID scores and their corresponding training iterations (d).
Recently, diffusion models have emerged as powerful deep generative modeling tools, showcasing remarkable performance in various applications, ranging from unconditional image generation [1, 2], conditional image generation [3, 4], image-to-image translation [5, 6, 7], text-to-image generation [8, 9, 10], inverse problem solving [11, 12, 13, 14], video generation [15, 16], and so on. These models employ a training process involving continuous injection of noise into training samples (“diffusion”), which are then utilized to generate new samples, such as images, by transforming random noise instances through a reverse diffusion process guided by the “score function [^2] The (Stein) score function of distribution $p_{t}$ at $\bm{x}_{t}$ is $\nabla_{\bm{x}_{t}}\log p_{t}(\bm{x}_{t})$. ” of the data distribution learned by the model. Moreover, recent work demonstrates that those diffusion models enjoy optimization stability and model reproducibility compared with other types of generative models [17]. However, diffusion models suffer from slow training and sampling despite their remarkable generative capabilities, which hinders their use in applications where real-time generation is desired [1, 2]. These drawbacks primarily arise from the necessity of tracking extensive forward and reverse diffusion trajectories, as well as managing a large model with numerous parameters across multiple timesteps (i.e., diffusion noise levels).
In this paper, we address these challenges based on two key observations: (i) there exists substantial parameter redundancy in current diffusion models, and (ii) they are trained inefficiently due to dissimilar gradients across different noise levels. Specifically, we find that training diffusion models require fewer parameters to accurately learn the score function at high noise levels, while larger parameters are needed at low noise levels. Furthermore, we also observe that when learning the score function, distinct shapes of distributions at different noise levels result in dissimilar gradients, which appear to slow down the training process driven by gradient descent.
Building on these insights, we propose a multi-stage framework with two key components: (1) a multi-decoder U-net architecture, and (2) a new partitioning algorithm to cluster timesteps (noise levels) into distinct stages. In terms of our new architecture, we design a multi-decoder U-Net that incorporates one universal encoder shared across all intervals and individual decoders tailored to each time stage; see fig. 1 (c) for an illustration. This approach combines the advantages of both universal and stage-specific architectures, which is much more efficient that using a single architecture for the entire training process [2, 1, 18] (fig. 1 (a)). Moreover, compared to previous approaches that completely separate architectures for each sub-interval [19, 20, 21, 22] (fig. 1 (b)), our approach can effectively mitigate interference between stages arising from disparate gradient effects, leading to improved efficiency. On the other hand, when it comes to partitioning the training stages of our network, we designed an algorithm aimed at grouping the timesteps. This is achieved by minimizing the functional distance within each cluster in the training objective and making use of the optimal denoiser formulation [18]. By integrating these two key components, our framework enables efficient allocation of computational resources (e.g., U-net parameters) and stage-tailored parameterization. In section 5, we showcase via extensive numerical experiments that our framework effectively improves both training and sampling efficiency. These experiments are performed on diverse benchmark datasets, demonstrating significant acceleration by using our framework when compared to three state-of-the-art (SOTA) diffusion model architectures.
Contributions.
As a summary, the major contributions of this work can be highlighted as follows:
  • Identifying two key sources of inefficiency. We identified two key sources that cause inefficiencies in training diffusion models across various time step: (i) a significant variation in the requirement of model capacity, and (ii) the gradient dissimilarity. As such, using a unified network cannot meet with the changing requirement at different time steps.
  • A new multi-stage framework. We introduced a new multi-stage architecture, illustrated in fig. 1 (c). We tackle these two sources of inefficiency by segmenting the time interval into multiple stages, where we employ customized multi-decoder U-net architectures that blends time-dependent models with a universally shared encoder.
  • Improved training and sampling efficiency. With comparable computational resources for unconditional image generation, we demonstrate that our multi-stage approach improves the Fréchet Inception Distance (FID) score for all SOTA methods. For example, on CIFAR-10 dataset [23], our method improves the FID for DPM-Solver [24] from 2.84 to 2.37, and it improves the FID for EDM [18] from 2.05 (our training result) to 1.96. Moreover, on the CelebA dataset [25], while maintaining a similar generation quality, our approach significantly reduces the required training FLOPS of EDM by 82% and the Latent Diffusion Model (LDM) [8] by 30%.
Organization.
In section 2, we provide preliminaries and an overview of related literatures. In section 3, we present our observations and analysis that motivated the proposed multistage framework, justifying its development. In section 4, we describe our proposed multistage framework for diffusion models, outlining the two core components. Finally, in section 5, we provide the results from our numerical experiments that validate the effectiveness of the proposed multistage approach.
2 Preliminaries & Related Work
In this section, we start by reviewing the basic fundamentals of diffusion models [1, 2, 18]. Subsequently, we delve into prior approaches aimed at improving the training and efficiency of diffusion models through the partitioning of the timestep interval. Lastly, we survey prior studies that significantly decrease the number of required sampling iterations.
Background on diffusion models. Let $\bm{x}_{0}\in\mathbb{R}^{n}$ denote a sample from the data distribution $p_{\text{data}}(\bm{x})$. Diffusion models operate within forward and reverse processes. The forward process progressively perturbs data $\bm{x}_{0}$ to a noisy version $\bm{x}_{t\in[0,1]}$ via corrupting with the Gaussian kernel. This process can be formulated as a stochastic differential equation (SDE)  [2] of the form $\text{d}\bm{x}_{t}=\bm{x}_{t}f(t)\text{d}t+g(t)\text{d}\bm{w}_{t}$, where $f(t)$ and $g(t)$ are the drift and diffusion coefficients, respectively, that correspond to a pre-defined noise schedule. $\bm{w}_{t}\in\mathbb{R}^{n}$ is the standard Wiener process. Under the forward SDE, the perturbation kernel is given by the conditional distribution defined as $p_{t}(\bm{x}_{t}|\bm{x}_{0})=\mathcal{N}(\bm{x}_{t};s_{t}\bm{x}_{0},s_{t}^{2}\sigma_{t}^{2}\textbf{I}),$ where
$s_{t}=\text{exp}(\int_{0}^{t}f(\xi)\text{d}\xi),\ \ \text{and}\ \sigma_{t}=\sqrt{\int_{0}^{t}\frac{g^{2}(\xi)}{s_{\xi}^{2}}\text{d}\xi}\>.$ (1)
The parameters $s_{t}$ and $\sigma_{t}$ are designed such that: (i) the data distribution is approximately estimated when $t=0$, and (ii) a nearly standard Gaussian distribution is obtained when $t=1$. The objective of diffusion models is to learn the corresponding reverse SDE, defined as $\text{d}\bm{x}=\left[f(t)\bm{x}_{t}-g^{2}(t)\nabla_{\bm{x}_{t}}\log p_{t}(\bm{x}_{t})\right]\text{d}t+g(t)\text{d}\bar{\bm{w}},$ where $\bar{\bm{w}}\in\mathbb{R}^{n}$ is the standard Wiener process running backward in time, and $\nabla_{\bm{x}_{t}}\log p_{t}(\bm{x}_{t})$ is the (Stein) score function. In practice, the score function is approximated using a neural network $\bm{\epsilon}_{\bm{\theta}}:\mathbb{R}^{n}\times[0,1]\rightarrow\mathbb{R}^{n}$ parameterized by $\bm{\theta}$, which can be trained by the denoising score matching technique [26] as
$\min\limits_{\bm{\theta}}\mathbb{E}\left[\omega(t)\|\bm{\epsilon}_{\bm{\theta}}(\bm{x}_{t},t)+s_{t}\sigma_{t}\nabla_{\bm{x}_{t}}\log p_{t}(\bm{x}_{t}|\bm{x}_{0})\|_{2}^{2}\right],$ (2)
which can also be written as $\min_{\bm{\theta}}\mathbb{E}[\omega(t)||\bm{\epsilon}_{\bm{\theta}}(\bm{x}_{t},t)-\bm{\epsilon}||^{2}]+C$, where the expectation is taken over $t\sim[0,1]$, $\bm{x}_{t}\sim p_{t}(\bm{x}_{t}|\bm{x}_{0})$, $\bm{x}_{0}\sim p_{\text{data}}(\bm{x})$, and $\bm{\epsilon}\sim\mathcal{N}(\textbf{0},\textbf{I})$. Here, $C$ is a constant independent of $\bm{\theta}$, and $\omega(t)$ is a scalar representing the weight of the loss as a function of $t$. In DDPM [1], it is simplified to $\omega(t)=1$. Once the parameterized score function $\bm{\epsilon}_{\bm{\theta}}$ is trained, it can be utilized to approximate the reverse-time SDE using numerical solvers such as Euler-Maruyama.
Timestep clustering methods. Diffusion models have demonstrated exceptional performance but face efficiency challenges in training and sampling. In response, several studies proposed to cluster the timestep range $t\in[0,1]$ into multiple intervals (e.g., $[0,t_{1}),[t_{1},t_{2}),\ldots,[t_{n},1]$). Notably, Choi et al. [19] reconfigured the loss weights for different intervals to enhance performance. Deja et al. [27] divide the entire process into a denoiesr and a generator based on their functionalities. Balaji et al. [28] introduced “expert denoisers”, which proposed using distinct architectures for different time intervals in text-to-image diffusion models. Go et al. [22] further improved the efficiency of these expert denoisers through parameter-efficient fine-tuning and data-free knowledge transfer. Lee et al. [21] designed separate architectures for each interval based on frequency characteristics. Moreover, Go et al. [20] treated different intervals as distinct tasks and employed multi-task learning strategies for diffusion model training, along with various timestep clustering methods.
Our approach distinguishes itself from the aforementioned methods in two key aspects. The first key component is our tailored U-net architecture using a unified encoder coupled with different decoders for different intervals. Previous models have either adopted a unified architecture, as seen in [19, 20], or employed separate architectures (referred to as expert denoisers) for each interval [28, 22, 21]. In comparison, our multistage architecture surpasses these methodologies, as demonstrated in section 5.3. Second, we developed a new timestep clustering method leveraging a general optimal denoiser (Prop. 1) that showcases superior performance (see section 5.4). In contrast, prior works rely on (i) a simple timestep-based clustering cost function [28, 20, 22, 21], (ii) Signal-to-Noise Ratio (SNR) based clustering [20], or (iii) gradients-based partitioning that uses task affinity scores [20].
Reducing the sampling iterations methods. Efforts to improve sampling efficiency of diffusion models have led to many recent advancements in SDE and Ordinary Differential Equation (ODE) samplers [2]. For instance, the Denoising Diffusion Implicit Model (DDIM) [29] formulates the forward diffusion as a non-Markovian process with a deterministic generative path, significantly reducing the number of function evaluations (NFE) required for sampling (from thousands to hundreds). Generalized DDIM (gDDIM) [30] further optimized DDIM by modifying the parameterization of the scoring network. Furthermore, the works in [24] and [31], termed the Diffusion Probabilistic Model solver (DPM-solver) and the Diffusion Exponential Integrator Sampler (DEIS), respectively, introduced fast higher-order solvers, employing exponential integrators that require 10s NFE for comparable generation quality. Moreover, the consistency model [32] introduced a novel training loss and parameterization, achieving high-quality generation with merely 1-2 NFE.
We remark that while the aforementioned methods are indirectly related to our work, our experiments in section 5.1 and section 5.2 show that our approach can be easily integrated into these techniques, further improving diffusion models’ overall training and sampling efficiency.
3 Identification of Key Sources of Inefficiency
Conventional diffusion model architectures, as exemplified by [2, 1, 18], treat the training of the diffusion model as a unified process across all timesteps. Recent researches, such as [19, 20, 21, 22], have highlighted the benefits of recognizing distinctions between different timesteps and the potential efficiency gains from treating them as separate tasks during the training process. However, our experimental results demonstrate that both unified and separate architectures suffer inefficiency for training diffusion models, where the inefficiency comes from (i) overparameterization, (ii) gradient dissimilarity, and (iii) overfitting.
3.1 Empirical Observations on the Key Sources of Inefficiency
(a) Analysis on interval $[0,t_{1})$
(b) Analysis on interval $[t_{2},1]$
Figure 2: Comparison between separate architecture and unified architecture w.r.t. the image generation quality in different intervals: (a) analysis on interval $[0,t_{1})$; and (b) analysis on interval $[t_{2},1]$. As illustrated on top of each figure, we only train separate architectures within specific intervals for the sampling process in both (a) and (b). For the remaining period of sampling, we use a well-trained diffusion model $(\bm{\epsilon}_{\bm{\theta}})^{[0,1]}_{4\times 10^{5}}$ to approximate the ground truth score function. As shown in the above figure of (a), e.g. for the separate architecture on interval 1, sampling utilizes trained model $(\bm{\epsilon}_{\bm{\theta}^{\prime}})^{[0,t_{1})}_{i}$ for interval 0 and well-trained model $(\bm{\epsilon}_{\bm{\theta}})^{[0,1]}_{4\times 10^{5}}$ for interval 1 and 2. Notably, for both $(\bm{\epsilon}_{\bm{\theta}})^{[0,1]}_{i}$ and $(\bm{\epsilon}_{\bm{\theta}})^{[0,1]}_{4\times 10^{5}}$ , we are using the model with 108M parameters. For separate architecture, the number in the parentheses represents the number of parameters of the model $(\bm{\epsilon}_{\bm{\theta}^{\prime}})^{[a,b]}_{i}$ For example, for separate architecture (169M) in (a), the model $(\bm{\epsilon}_{\bm{\theta}^{\prime}})^{[0,t_{1})}_{i}$ has 169M parameters for $\bm{\theta}^{\prime}$. The bottom figures in (a-b) illustrate the FID of the generation from each architecture under different training iterations.
To illustrate the inefficiency in each interval, we isolate the interval by using a separate architecture from the rest.
Experiment setup.
In our experiments, we consider three-stage training and divide the time steps into three intervals: $[0,t_{1}),[t_{1},t_{2}),[t_{2},1]$ [^3] Details for interval clustering can be found in section 5.. Let $(\bm{\epsilon}_{\bm{\theta}})^{[a,b]}_{i},0\leq a<b\leq 1$ denote a U-Net architecture with parameter $\bm{\theta}$ trained with $i$ iterations and fed with data pairs $(\bm{x}_{t},t)$, where $t\in[a,b]$. We then train models using two different strategies: a unified architecture with 108M network parameters for all intervals, i.e., $(\bm{\epsilon}_{\bm{\theta}})^{[0,1]}_{i}$, and separate architectures with varying network parameters (e.g., 47M, 108M, 169M) for each interval; e.g., $(\bm{\epsilon}_{\bm{\theta}})^{[0,t_{1})}_{i}$ for the interval $[0,t_{1})$, etc. It is worth noting that, apart from the differences in network parameters, we utilize the same network architecture (e.g., U-Net) for both the unified and the separate approaches. We assessed the training progress of each model by evaluating image generation quality at different training iterations. Notably, because some of the models are only trained on one interval, we need to provide a ground truth score for the other intervals. In fig. 2, the sampling process is shown top and the experimental results are shown bottom.
Inefficiency in unified architectures.
From fig. 2, we observe the following:
  • Overparameterization and underfitting emerge simultaneously for unified architectures. In fig. 2(a), we observe that increasing the number of parameters in Interval 0 can improve the image generation quality (as indicated by a lower FID score). In contrast, fig. 2(b) reveals that increasing the number of parameters in Interval 2 has minimal impact on the quality of image generation. This implies that using a unified architecture will result in underfitting in Interval 0 and overparameterization in Interval 2. The current unified architecture’s parameter redundancy leaves significant room for improving its efficiency. To optimize the computational usage, we should allocate more parameters to Interval 0 while allocating fewer parameters to Interval 2.
  • Gradient dissimilarity hinders training for unified architecture. Quantitative results from [20] demonstrate dissimilarity in gradients caused by differences within each interval. This can also be observed from our results based upon both fig. 2(a) and fig. 2(b). For the unified and separate architectures using the same number of parameters (108M), separate architecture achieves a significantly lower FID with the same training iterations, implying that dissimilar gradients among intervals may hinder training when using a unified architecture. Here, the only difference between training separate and unified architectures is that the batch gradient for unified architecture is calculated based on all timesteps while the batch gradien2 for separate architecture is calculated only from a specific interval.
Inefficiency in existing separate architectures.
Although separated architecture [19, 21, 22] better allocates computational resources for each interval, it suffers from overfitting. This can be illustrated based upon training separate architectures (169M) and (108M) in Interval 0 shown in fig. 2(a), where increasing the number of parameters will lead to overfitting. This also happens in Interval 2, when we compare all separate architectures in fig. 2(b). In comparison, the unified networks with 108M parameters are less prone to overfitting for both Interval 0 and Interval 2. This suggests that we can reduce overfitting by training shared weights across different intervals together.
3.2 Tackling the Inefficiency via Multistage U-Net Architectures
In a unified architecture applied across all timesteps, there is often a dual challenge: requirements for more parameters (169M) in the interval $[0,t_{1})$ but fewer parameters (47M) in the interval $[t_{2},1]$. This issue is compounded by the gradient dissimilarity across different timesteps, which can impede effective training. Alternatively, employing separate architectures for different intervals might lead to overfitting and a lack of robust early stopping mechanisms. To address these challenges, our proposed multistage architecture in section 4 combines shared parameters to reduce overfitting with interval-specific parameters to mitigate the impact of gradient dissimilarity. This tailored approach for each interval ensures improved adaptability. Furthermore, we conduct an in-depth ablation study in section 5.3 to showcase the effectiveness of our multi-stage architecture over the existing models.
4 Proposed Multistage Framework
In this section, we introduce our new multistage framework (as illustrated in fig. 1 (c)). Specifically, we first introduce the multi-stage U-Net architecture design in section 4.1, following a new clustering method for choosing the optimal interval to partition the entire timestep $[0,1]$ into intervals in section 4.2. , and discuss the rationale of the proposed architecture in section 4.3.
4.1 Proposed Multi-stage U-Net Architectures
As discussed in section 3, most existing diffusion models either employ a unified architecture across all intervals [1, 2, 18] to share features for all timesteps, or use completely separate architectures for different timestep intervals [21, 22, 28] where the goal is to take advantage of the benign properties within different intervals.
To harness the advantages of both unified and separate architectures employed in prior studies, we introduce a multistage U-Net architecture, as illustrated in Figure 1(c). Specifically, we partition the entire timestep $[0,1]$ into several intervals, e.g., three intervals $[0,t_{1}),[t_{1},t_{2}),[t_{2},1]$ in fig. 1. For the architecture, we introduce:
  • One shared encoder across all time intervals. For each timestep interval, we implement a shared encoder architecture (plotted in blue in fig. 1 (c)), which is similar to the architecture employed in the original U-Net framework [33]. Unlike separate architecture, the shared encoder provide shared information across all timesteps, preventing models from overfitting (see section 5.3 for a discussion).
  • Separate decoders for different time intervals. Motivated by the multi-head structure introduced in the Mask Region-based Convolutional Neural Networks (Mask-RCNN) method [34], we propose to use multiple distinct decoders (plotted in colors for different intervals in fig. 1(c)), where each decoder is tailored to a specific timestep interval. The architecture of each decoder closely resembles the one utilized in [2], with deliberate adjustments made to the embedding dimensions to optimize performance.
As we observe, the primary difference in the architecture resides within the decoder structure. Intuitively, we use a decoder with fewer number of parameters for intervals closer to the noise, because the learning task is easier. We use a decoder with a larger number of parameters for intervals closer to the image.
4.2 Optimal Denoiser-based Timestep Clustering
Next, we discuss how we principally choose the interval partition time points in practice. For simplicity, we focus on the case where we partition the time $[0,1]$ into three intervals $[0,t_{1}),[t_{1},t_{2}),[t_{2},1]$, and we develop a timestep clustering method to choose the optimal $t_{1}$ and $t_{2}$. Of course, our method can be generalized to multi-stage networks with arbitrary interval numbers. However, in practice, we find that the choice of three intervals strikes a good balance between effectiveness and complexity; see our ablation study in section B.5.
To partition the time interval, we employ the optimal denoiser established in Proposition 1.
Algorithm 1 Optimal Denoiser based Timestep Clustering
1:Input: Total samples $K$, optimal denoiser function $\bm{\epsilon}^{*}_{\bm{\theta}}(\bm{x},t)$, thresholds $\alpha$, $\eta$, dataset $p_{\text{data}}$, $\mathcal{S}_{0}=\mathcal{S}_{1}=\emptyset$
2:Output: Timesteps $t_{1}$, $t_{2}$
3:for $k\in\{1,\dots,K\}$ do
4:    $\bm{y}_{k}\sim p_{\text{data}},\bm{\epsilon}_{k}\sim\mathcal{N}(0,\mathbf{I}),t_{k}\sim[0,1]$
5:    $\mathcal{S}^{k}_{0}\leftarrow\mathcal{D}(\bm{\epsilon}^{*}_{t_{k}},\bm{\epsilon}^{*}_{0},\bm{y}_{k},\bm{\epsilon}_{k}),\mathcal{S}^{k}_{1}\leftarrow\mathcal{D}(\bm{\epsilon}^{*}_{t_{k}},\bm{\epsilon}^{*}_{1},\bm{y}_{k},\bm{\epsilon}_{k})$
6:    $\mathcal{S}_{0}\leftarrow\mathcal{S}_{0}\cup\{\left(t_{k},\mathcal{S}^{k}_{0}\right)\},\mathcal{S}_{1}\leftarrow\mathcal{S}_{1}\cup\{\left(t_{k},\mathcal{S}^{k}_{1}\right)\}$
7:end for
8:$t_{1}=\operatorname*{\text{arg~{}max}}\limits_{\tau}\left\{\tau~{}\Big{|}~{}\frac{\sum_{(t_{k},\mathcal{S}^{k}_{0})\in\mathcal{S}_{0}}\left[\mathcal{S}^{k}_{0}\cdot\mathds{1}(t_{k}\leq\tau)\right]}{\sum_{(t_{k},\mathcal{S}^{k}_{0})\in\mathcal{S}_{0}}\left[\mathds{1}(t_{k}\leq\tau)\right]}\geq\alpha\right\}$
9:$t_{2}=\operatorname*{\text{arg~{}min}}\limits_{\tau}\left\{\tau~{}\Big{|}~{}\frac{\sum_{(t_{k},\mathcal{S}^{k}_{1})\in\mathcal{S}_{0}}\left[\mathcal{S}^{k}_{1}\cdot\mathds{1}(t_{k}\geq\tau)\right]}{\sum_{(t_{k},\mathcal{S}^{k}_{1})\in\mathcal{S}_{0}}\left[\mathds{1}(t_{k}\geq\tau)\right]}\geq\alpha\right\}$
Proposition 1.
Suppose we train a diffusion model denoiser function $\bm{\epsilon}_{\bm{\theta}}(\bm{x},t)$ with parameters $\bm{\theta}$ using dataset $\left\{\bm{y}_{i}\in\mathbb{R}^{n}\right\}_{i=1}^{N}$ by
$\min_{\bm{\theta}}\mathcal{L}(\bm{\epsilon}_{\bm{\theta}};t)=\mathbb{E}_{\bm{x}_{0},\bm{x}_{t},\epsilon}[||\bm{\epsilon}-\bm{\epsilon}_{\bm{\theta}}(\bm{x}_{t},t)||^{2}],$ (3)
where $\bm{x}_{0}\sim p_{\text{data}}(\bm{x})=\frac{1}{N}\sum_{i=1}^{N}\delta(\bm{x}-\bm{y}_{i})$, $\bm{\epsilon}\sim\mathcal{N}(0,\textbf{I})$, and $\bm{x}_{t}\sim p_{t}(\bm{x}_{t}|\bm{x}_{0})=\mathcal{N}(\bm{x}_{t};s_{t}\bm{x}_{0},s_{t}^{2}\sigma_{t}^{2}\textbf{I})$ with perturbation parameters $s_{t}$ and $\sigma_{t}$ defined in eq. 1. Then, the optimal denoiser at $t$, defined as $\bm{\epsilon}^{*}_{\bm{\theta}}(\bm{x};t)=\operatorname*{\text{arg~{}min}}_{\bm{\epsilon}_{\bm{\theta}}}\mathcal{L}(\bm{\epsilon}_{\bm{\theta}};t)$, is given by
$\bm{\epsilon}^{*}_{\bm{\theta}}(\bm{x};t)=\frac{1}{s_{t}\sigma_{t}}\left[\bm{x}-s_{t}\frac{\sum_{i=1}^{N}\mathcal{N}(\bm{x};s_{t}\bm{y}_{i},s_{t}^{2}\sigma_{t}^{2}\textbf{I})\bm{y}_{i}}{\sum_{i=1}^{N}\mathcal{N}(\bm{x};s_{t}\bm{y}_{i},s_{t}^{2}\sigma_{t}^{2}\textbf{I})}\right]\>.$ (4)
The proof is provided in appendix A, and the result can be generalized from recent work of Karras et al. [18], extending from a specific kernel $p_{t}(\bm{x}_{t}|\bm{x}_{0})=\mathcal{N}(\bm{x}_{t};\bm{x}_{0},\sigma_{t}^{2}\textbf{I})$ to encompassing a broader scope of noise perturbation kernels, given by $p_{t}(\bm{x}_{t}|\bm{x}_{0})=\mathcal{N}(\bm{x}_{t};s_{t}\bm{x}_{0},s_{t}^{2}\sigma_{t}^{2}\textbf{I})$. For brevity, we simplify the notation of the optimal denoiser $\bm{\epsilon}^{*}_{\theta}(\bm{x},t)$ in Proposition 1 as $\bm{\epsilon}^{*}_{t}(\bm{x})$.
To obtain the optimal interval, our rationale is to homogenize the regression task as much as possible within each individual time interval. To achieve this goal, given sampled $\bm{x}_{0}$, $\bm{\epsilon}$, we define the function distance of the optimal denoiser at any given timestep $t_{a}$, $t_{b}$ as:
$\mathcal{D}(\bm{\epsilon}^{*}_{t_{a}},\bm{\epsilon}^{*}_{t_{b}},\bm{x}_{0},\bm{\epsilon})=\frac{1}{n}\sum_{i=1}^{n}\mathds{1}(|\bm{\epsilon}^{*}_{t_{a}}(\bm{x}_{t_{a}})-\bm{\epsilon}^{*}_{t_{b}}(\bm{x}_{t_{b}})|_{i}\leq\eta)\>,$
where $\mathds{1}(\cdot)$ is the indicator function, $\eta$ is a pre-specified threshold, $\bm{x}_{t_{a}}=s_{t_{a}}\bm{x}_{0}+s_{t_{a}}\sigma_{t_{a}}\bm{\epsilon}$, and $\bm{x}_{t_{b}}=s_{t_{b}}\bm{x}_{0}+s_{t_{b}}\sigma_{t_{b}}\bm{\epsilon}$. Consequently, we define the functional similarity of the optimal denoiser at timesteps $t_{a}$ and $t_{b}$ as:
$\mathcal{S}(\bm{\epsilon}^{*}_{t_{a}},\bm{\epsilon}^{*}_{t_{b}})=\mathbb{E}_{\bm{x}_{0}\sim p_{\text{data}}}\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(0,\textbf{I})}[\mathcal{D}(\bm{\epsilon}^{*}_{t_{a}},\bm{\epsilon}^{*}_{t_{b}},\bm{x}_{0},\bm{\epsilon})]\>.$ (5)
Based upon the definition, we design the following optimization problem to find the largest $t_{1}$ and smallest $t_{2}$ as:
$t_{1}$ $\leftarrow\operatorname*{\text{arg~{}max}}\limits_{\tau}\left\{\tau\Big{|}\mathbb{E}_{t\sim[0,\tau)}[\mathcal{S}(\bm{\epsilon}^{*}_{t},\bm{\epsilon}^{*}_{0})]\geq\alpha\right\},$ (6)
$t_{2}$ $\leftarrow\operatorname*{\text{arg~{}min}}\limits_{\tau}\left\{\tau\Big{|}\mathbb{E}_{t\sim[\tau,1]}[\mathcal{S}(\bm{\epsilon}^{*}_{t},\bm{\epsilon}^{*}_{1})]\geq\alpha\right\},$ (7)
such that the average functional similarity of $\bm{\epsilon}^{*}_{t}$ (resp. $\bm{\epsilon}^{*}_{t}$) to $\bm{\epsilon}^{*}_{0}$ (resp. $\bm{\epsilon}^{*}_{1}$) in $[0,t_{1})$ (resp. $[t_{2},1]$) is larger than or equal to a pre-defined threshold $\alpha$. As the above optimization problems are intractable, we propose the procedure outlined in algorithm 1 to obtain an approximate solution. In particular, the algorithm samples $K$ pairs $(\bm{y}_{k},\bm{\epsilon}_{k},t_{k}),k\in\{1,\ldots,K\}$ to calculate the distances $\mathcal{D}(\bm{\epsilon}^{*}_{t_{k}},\bm{\epsilon}^{*}_{0},\bm{y}_{k},\bm{\epsilon}_{k})$ and $\mathcal{D}(\bm{\epsilon}^{*}_{t_{k}},\bm{\epsilon}^{*}_{1},\bm{y}_{k},\bm{\epsilon}_{k})$ (step 6). Based on those distances, we solve the optimization problems defined in the lines 8 and 9 of algorithm 1 to obtain $t_{1}$ and $t_{2}$.
4.3 Rationales for the proposed architecture
Finally, we summarize the rationales of our proposed architecture based on the empirical observations in section 3 and previous works.
Rationales for the shared encoder. (i) Prevent overfitting: If we treat training separate stages of diffusion models as multi-task learning, [35] suggests that shared parameters across tasks in multi-task learning can mitigate overfitting. (ii) Maintain consistency in h-space. The output of the encoder for UNet is named as h-space [36], which has properties for semantic manipulation such as homogeneity, linearity, robustness, and consistency across timesteps. Therefore, sharing the encoder can maintain better consistency of the h-space among all timesteps compared to separate encoders.
Rationales for network parameter design. First, we provide some intuitions as to why the learning task at $t=0$ is harder than that at $t=1$. Suppose diffusion models could converge to the optimal denoiser $\bm{\epsilon}_{\bm{\theta}}^{*}(\bm{x}_{t},t)$ given in eq. 4. Based on this, we observe: (i) when $t\to 0$, we have $\bm{x}_{t\to 0}=\bm{x}_{0}$, $\bm{\epsilon}_{\bm{\theta}}^{*}|_{t\to 0}=\bm{\epsilon}$, so that $\bm{\epsilon}^{*}_{\bm{\theta}}$ is a complicated mapping from training data distribution $p_{\text{data}}(\bm{x})$ to the gaussian distribution; (ii) when $t\to 1$, we have $\bm{x}_{t\to 1}=\bm{\epsilon}$, $\bm{\epsilon}_{\bm{\theta}}^{*}|_{t\to 1}=\bm{\epsilon}$, so $\bm{\epsilon}^{*}_{\bm{\theta}}$ is an identity mapping. The two extreme cases reveal that the identity mapping for $t\to 1$ (close to noise) is easier to learn than that for $t\to 0$. Second, our choice of network parameters across different stages share similar spirits with the recent work [37], which employs a high dimensional subspace as $t\to 0$ and gradually decreases the dimensionality of each subspace until $t=1$.
5 Experiments
For this section, we start with providing the basic experimental setups. Followed by this, in section 5.1, we demonstrate the improved image generation quality of our multistage architecture over the state-of-the art models, under comparable training and sampling computations. On the other hand, in 5.2, we demonstrate the improved training and sampling efficiency of our methods over existing methods, when compared with similar image generation quality. Finally, in 5.3 and 5.4, we conduct comprehensive ablation studies on timestep clustering methods and multistage architectures, illustrating the superiority of the choices in our method.
Settings of the timestep clustering method.
We implement a Variance Preserving (VP) [1, 18] perturbation kernel for clustering the time interval. In particular, we use $s_{t}=\sqrt{e^{\frac{1}{2}\beta_{d}t^{2}+\beta_{\text{min}}-1}},\sigma_{t}=1/\sqrt{e^{\frac{1}{2}\beta_{d}t^{2}+\beta_{\text{min}}}},t\in[\epsilon_{t},1],\beta_{d}=19.9,\beta_{\text{min}}=0.1$, and $,\epsilon_{t}=10^{-3}$. To obtain $t_{1}$ and $t_{2}$, we utilize the CIFAR-10 [23] with dataset size $N=5\times 10^{4}$. Specifically, Algorithm 1 is configured with $\eta=\frac{2}{256},\alpha=0.9,$ and $K=5\times 10^{4}$. This choice of $\eta$ ensures the similarity of measurements in the RGB space. We divide $t\in[0,1]$ into $10^{4}$ discrete time steps. Consequently, we employ a grid search to determine optimal values for $t_{1}$ and $t_{2}$ using the procedure outlined in Algorithm 1. For the CelebA dataset [25], we utilize the same $t_{1}$ and $t_{2}$. And for Variance Exploding (VE) perturbation kernel, we calculate an equivalent $\sigma_{1}$ and $\sigma_{2}$. [^4] VE utilizes $\sigma$ instead of $t$ to represent the timestep. We choose $\sigma_{1}$ and $\sigma_{2}$ such that $\text{SNR}_{\text{VE}}(\sigma_{1})=\text{SNR}_{\text{VP}}(t_{1})$ and $\text{SNR}_{\text{VE}}(\sigma_{2})=\text{SNR}_{\text{VP}}(t_{2})$. Further details are provided in the appendices. Results of Section 5.4 verify that these choices of $t_{1}$ and $t_{2}$ return the best performance when compared to other clustering method.
Multistage architectures.
Our multistage architecture, inspired by the U-Net model [33] used in DDPM++ [1, 18, 2], is modified for interval-specific channel dimensions. The proposed architecture is adopted to three diffusion models: DPM-Solver [24], Elucidating Diffusion Model (EDM) [18], and Latent Diffusion Model (LDM) [8]. In particular, for the cases of DPM-Solver and EDM, the encoder’s channel dimensions are standardized at 128, while the decoders are configured with 192, 128, and 16 channels for intervals $[0,t_{1})$, $[t_{1},t_{2})$, and $[t_{2},1]$, respectively. In the LDM case, we use 224 channels across the encoder for all intervals whereas the decoders are configured with 256, 192, and 128 channels for the respective intervals. To decide the specific number of parameters for each decoder, we apply ablation studies in the appendices.
Datasets, Evaluation Metrics, & Baselines.
We use CIFAR-10 (32 $\times$ 32), CelebA (32 $\times$ 32), and CelebA (256 $\times$ 256) datasets for our experiments. To evaluate the performance of our multistage diffusion model in terms of the generation quality, we use the standard Fréchet Inception Distance (FID) metric [38], and the required Number of Function Evaluations (NFEs). We assess the sampling efficiency using giga-floating point operations (GFLOPs) per function evaluation. For both separate architecture and our multistage architecture, equivalent GFLOPs are computed as a weighted summation (based on the ODE Sampler steps within each interval) of GFLOPs for each interval. Training efficiency is evaluated using total training iterations multiplied by the floating point operations per function evaluation, [^5] Here we simplify it by ignoring the FLOPs for backward propagation, which is approximately proportional to FLOPs of forward evaluation. measured by peta-floating point operations (PFLOPs). We choose DDPM [1], Score SDE [2], Poisson Flow Generative Models (PFGM) [39], DDIM [29], gDDIM [30], DEIS [31], DPM-solver [24], and EDM [18] as the baseline.
5.1 Image Generation Quality Results
In this subsection, we demonstrate the effectiveness of our approach by comparing the image generation quality (measured by FID) with comparable training and sampling computations (measured by NFE). Specifically, table 1 presents FID scores to measure the sampling quality, and NFE to measure the number of sampling iterations required using the CIFAR-10 dataset. Our method is compared to 8 baselines. As observed, our multistage DPM-Solver outperforms DPM-Solver in terms of the reported FID values while requiring similar training iterations (both are $4.5\times 10^{5}$) and model GFLOPs (18.65 for multistage DPM-Solver versus 17.65 for DPM-Solver). A similar observation holds when we compare our multistage EDM and the vanilla EDM, where we reduce FID from 2.05 to 1.96 by using the multi-stage architecture. Remarkably, utilizing only 20 NFE, our Multistage DPM-Solver returns the same FID score as the one reported for the PFGM method, which requires 147 NFEs. These results also highlight the adaptability of our framework to higher-order ODE solvers; see the 8th and last row of table 1.
METHOD NFE ($\downarrow$) FID ($\downarrow$)
DDPM 1000 3.17
Score SDE 2000 2.20
PFGM 147 2.35
DDIM 100 4.16
gDDIM 20 2.97
DEIS 20 2.86
DPM-solver 20 2.73
Multistage DPM-solver (Ours) 20 2.35
EDM 35 2.05
Multistage EDM (Ours) 35 1.96
Table 1: Sampling quality on CIFAR-10 dataset.
5.2 Training & Sampling Efficiency Results
Dataset/Method Training Iterations ($\downarrow$) GFLOPs ($\downarrow$) Total Training PFLOPs ($\downarrow$) FID ($\downarrow$)
CIFAR-10 $32\times 32$
DPM-Solver [18] $4.5\times 10^{5}$ 17.65 7.94 2.73
[…]
CelebA $32\times 32$
EDM [18] $5.7\times 10^{5}$ 17.65 10.06 1.55
[…]
CelebA $256\times 256$
LDM [8] $4.9\times 10^{5}$ 88.39 43.31 8.29
[…]
Table 2: Training and sampling efficiency on more datasets.
In this subsection, we further demonstrate the superiority of our method by comparing the training and sampling efficiency under comparable image generation quality. Specifically, in table 2, we present the number of training iterations, GFLOPs, and total training PFLOPs of our approach, DPM-solver, EDM, and LDM using CIFAR-10 and CelebA datasets. Using the CIFAR-10 dataset, our multistage DPM-solver achieves similar FID scores (2.71 vs 2.73) while requiring nearly half the training iterations when compared to the vanilla DPM-solver. For the case of EDM (resp. LDM), our approach returns an FID score of 1.44 (resp. 8.29), requiring $1.4\times 10^{5}$ ($3.2\times 10^{5}$) less iterations when compared to vanilla DPM-solver (resp. LDM). For the cases of DPM-solver and EDM, we can achieve a substantial reduction of training iterations, which is demonstrated by a marginal increase in the number of GFLOPs. For the LDM case, we also achieve a significant reduction of both training iterations and GFLOPs. These promising results highlight the significantly improved computational efficiency achieved by using the proposed multistage framework.
5.3 Comparison of Different Architectures
In section 3, we highlighted the limitations of both unified and separate diffusion model architectures in terms of training efficiency (see fig. 2). In this part, we further illustrate these limitations through extensive experiments as shown in table 3. Here, we use the U-Net architecture, trained on the CIFAR-10 dataset, and utilize the DPM-Solver for sampling. For the unified case, we use a single U-Net model with 128 channels. For the separate case, three distinct U-Nets with 128 channels are used. For improved performance of the separate architecture, we implement two techniques: early stopping (ES) and tailored parameters (TP) to tackle the overfitting and parameter inefficiency discussed in section 3. Under ES, the criteria is to stop training prior to overfitting. For TP, the three U-Nets are configured with 192, 128, and 16 channels decoders for Intervals 0, 1, and 2, respectively.
Method GFLOPs FID ($\downarrow$)
Unified 17.65 2.73
Separate 17.65 2.87
Separate (+ ES) 17.65 2.80
Separate (+ TP) 18.65 2.68
Separate (+ ES, TP) 18.65 2.52
Multistage (Ours) 18.65 2.35
Table 3: Sampling quality on CIFAR-10 dataset for different diffusion model architectures. Here, ES denotes early stopping, and TP denotes tailored parameters.
Our comparison and analysis in table 3 reveal notable insights of our network design. Comparisons between the 2nd and 3rd rows (and between the 4th and 5th rows) on the separate architectures indicate that early stopping effectively mitigates overfitting and enhances generation quality. When comparing the 2nd and 4th rows (as well as the 3rd and 5th rows) on the separate architectures, we observe that optimizing parameter usage can achieve a significant decrease in FID under comparable GFLOPs. Most importantly, our multistage architecture, as shown in the 6th row, benefits from both unified and separate architectures, achieving the best FID (2.35, compared to 2.73 and 2.52). Comparing the 2nd row and the 4th row, the shared encoder not only prevents overfitting but also improves the convergence of the diffusion model as reported by the FID scores.
5.4 Comparison of Timestep Clustering Methods
Clustering Method $t_{1}$ $t_{2}$ FID ($\downarrow$)
Timestep [20, 21] 0.330 0.670 3.12
SNR [20] 0.348 0.709 2.72
Gradient [20] 0.360 0.660 2.75
Optimal Denoiser (Ours) 0.442 0.631 2.35
Table 4: Sampling quality on CIFAR-10 dataset for different clustering methods.
As previously stated in section 2, various timestep clustering methods are proposed including timestep-based, SNR-based, and gradient-based clustering approaches [20, 21]. In this subsection, we conduct an experiment to demonstrate the superiority of our clustering method compared to previous arts. Specifically, we apply the clustering methods in [20, 21] to partition the interval along with our proposed multistage UNet architecture. The computed intervals are shown in the table 4. We use the multistage DPM-Solver with these different intervals trained on the CIFAR-10 dataset. As observed, our optimal denoiser-based clustering method achieves the highest FID score, consistently outperforming all other clustering methods.
6 Conclusion & Future Work
In this paper, we introduced a novel multi-stage framework for diffusion models (DM) to improve the training and sampling efficiency. We proposed an algorithm that divides the timestep into several stages. Based on these stages, we designed a stage-specific multi-decoder U-net architecture and a shared encoder across all stages. We conducted thorough numerical experiments with several SOTA diffusion model frameworks and confirmed the effectiveness of our strategy using small scale and large scale datasets.
In future research, it would be interesting to expand our multi-stage approach beyond unconditional diffusion models by considering conditional DMs and DM-based inverse problems solvers. Our experiment in section 5.2 demonstrate that training latent diffusion models within our multi-stage framework requires only 30% of the computational effort needed for training standard latent diffusion models on the CelebA dataset. Thus, employing a multi-stage strategy could significantly reduce the computational demands for training large-scale stable diffusion models, such as those described in [8], which typically requires significant computations.
Acknowledgment
HJZ, YFL, and QQ acknowledge support from NSF CAREER CCF-2143904, NSF CCF-2212066, NSF CCF-2212326, NSF IIS 2312842, ONR N00014-22-1-2529, an AWS AI Award, and a gift grant from KLA. QQ also acknowledges the support from MICDE Catalyst Grant. IA and SR acknowledge the support from the NSF grants CCF-2212065 and CCF-2212066. Results presented in this paper were obtained using CloudBank, which is supported by the NSF under Award #1925001, and the authors acknowledge efficient cloud management framework SkyPilot [40] for computing. The authors acknowledge valuable discussions with Prof. Jeffrey Fessler (U. Michigan), Prof. Liyue Shen (U. Michigan), and Prof. Rongrong Wang (MSU).
Appendix A Proof of Proposition 1
This section presents the proof of Prop. 1 in section 4.
Given dataset $\left\{\bm{y}_{i}\right\}_{i=1}^{N}$ with $N$ images, we approximate the original dataset distribution as $p_{\text{data}}$ as multi-Dirac distribution, $p_{\text{data}}(\bm{x})=\frac{1}{N}\sum_{i=1}^{N}\delta(\bm{x}-\bm{y}_{i})$. Then, the distribution of the perturbed image $\bm{x}$ at random timestep $t$ can be calculated as:
$p_{t}(\bm{x})$ $=\int_{\mathbb{R}^{d}}p_{t}(\bm{x}|\bm{x}_{0})p_{\text{data}}(\bm{x}_{0})\text{d}\bm{x}_{0}$ (8)
$=\int_{\mathbb{R}^{d}}p_{\text{data}}(\bm{x}_{0})\mathcal{N}(\bm{x};s_{t}\bm{x}_{0},s_{t}^{2}\sigma_{t}^{2}\textbf{I})\text{d}\bm{x}_{0}$ (9)
$=\int_{\mathbb{R}^{d}}\frac{1}{N}\sum_{i=1}^{N}\delta(\bm{x}_{0}-\bm{y}_{i})\mathcal{N}(\bm{x};s_{t}\bm{x}_{0},s_{t}^{2}\sigma_{t}^{2}\textbf{I})\text{d}\bm{x}_{0}$ (10)
$=\frac{1}{N}\sum_{i=1}^{N}\int_{\mathbb{R}^{d}}\delta(\bm{x}_{0}-\bm{y}_{i})\mathcal{N}(\bm{x};s_{t}\bm{x}_{0},s_{t}^{2}\sigma_{t}^{2}\textbf{I})\text{d}\bm{x}_{0}$ (11)
$=\frac{1}{N}\sum_{i=1}^{N}\mathcal{N}(\bm{x};s_{t}\bm{y}_{i},s_{t}^{2}\sigma_{t}^{2}\textbf{I})\>.$ (12)
Let the noise prediction loss, which is generally used across various diffusion models, be
$\mathcal{L}(\bm{\epsilon}_{\bm{\theta}};t)$ $=\mathbb{E}_{\bm{x}\sim p_{t}(\bm{x})}[|\bm{\epsilon}-\bm{\epsilon}_{\bm{\theta}}(\bm{x},t)||^{2}]=\int_{\mathbb{R}_{d}}\frac{1}{N}\sum_{i=1}^{N}\mathcal{N}(\bm{x};s_{t}\bm{y}_{i},s_{t}^{2}\sigma_{t}^{2}\textbf{I})||\bm{\epsilon}-\bm{\epsilon}_{\bm{\theta}}(\bm{x},t)||^{2}\text{d}\bm{x}\>,$ (13)
where $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\textbf{I})$ is defined follow the perturbation kernel $p_{t}(\bm{x}|\bm{x}_{0})=\mathcal{N}(\bm{x};s_{t}\bm{x}_{0},s_{t}^{2}\sigma_{t}^{2}\textbf{I})$, and
$\bm{x}=s_{t}\bm{y}_{i}+s_{t}\sigma_{t}\bm{\epsilon}\quad\Rightarrow\quad\bm{\epsilon}=\frac{\bm{x}-s_{t}\bm{y}_{i}}{s_{t}\sigma_{t}}\>.$ (14)
Here, $\bm{\epsilon}_{\bm{\theta}}$ is a "denoiser" network for learning the noise $\bm{\epsilon}$. Plugging (14) into (13), the loss can be reparameterized as:
$\mathcal{L}(\bm{\epsilon}_{\bm{\theta}};t)=\int_{\mathbb{R}_{d}}\mathcal{L}(\bm{\epsilon}_{\bm{\theta}};\bm{x},t)\text{d}\bm{x}\>,$ (15)
where
$\mathcal{L}(\bm{\epsilon}_{\bm{\theta}};\bm{x},t)=\frac{1}{N}\sum_{i=1}^{N}\mathcal{N}(\bm{x};s_{t}\bm{y}_{i},s_{t}^{2}\sigma_{t}^{2}\textbf{I})||\bm{\epsilon}_{\bm{\theta}}(\bm{x},t)-\frac{\bm{x}-s_{t}\bm{y}_{i}}{s_{t}\sigma_{t}}||^{2}\>.$
Eq.(15) means that we can minimize $\mathcal{L}(\bm{\epsilon}_{\bm{\theta}};t)$ by minimizing $\mathcal{L}(\bm{\epsilon}_{\bm{\theta}};\bm{x},t)$ for each $\bm{x}$. As such, we find the "optimal denoiser" $\bm{\epsilon}^{*}_{\bm{\theta}}$ that minimize the $\mathcal{L}(\bm{\epsilon}_{\bm{\theta}};\bm{x},t)$, for every given $\bm{x}$ and $t$, as:
$\bm{\epsilon}^{*}_{\bm{\theta}}(\bm{x};t)=\text{arg}\ \text{min}_{\bm{\epsilon}_{\bm{\theta}}(\bm{x};t)}\mathcal{L}(\bm{\epsilon}_{\bm{\theta}};\bm{x},t)\>.$ (16)
The above equation is a convex optimization problem with respect to $\bm{\epsilon}_{\bm{\theta}}$ by which the solution can be obtained by setting the gradient of $\mathcal{L}(\bm{\epsilon}_{\bm{\theta}};\bm{x},t)$ w.r.t $\bm{\epsilon}_{\bm{\theta}}(\bm{x};t)$ to zero. Under the assumption that $\bm{\epsilon}_{\bm{\theta}}$ has an infinite model capacity, and can approximate any continuous function to an arbitrary level of accuracy based on the Universal Approximation Theorem [41], then
$\nabla_{\bm{\epsilon}_{\bm{\theta}}(\bm{x};t)}[\mathcal{L}(\bm{\epsilon}_{\bm{\theta}};\bm{x},t)]=0$ (17)
$\Rightarrow$ $\frac{1}{N}\sum_{i=1}^{N}\mathcal{N}(\bm{x};s_{t}\bm{y}_{i},s_{t}^{2}\sigma_{t}^{2}\textbf{I})[\bm{\epsilon}^{*}_{\bm{\theta}}(\bm{x};t)-\frac{\bm{x}-s_{t}\bm{y}_{i}}{s_{t}\sigma_{t}}]=0$ (18)
$\Rightarrow$ $\bm{\epsilon}^{*}_{\bm{\theta}}(\bm{x};t)=\frac{1}{s_{t}\sigma_{t}}[\bm{x}-s_{t}\frac{\sum_{i=1}^{N}\mathcal{N}(\bm{x};s_{t}\bm{y}_{i},s_{t}^{2}\sigma_{t}^{2}\textbf{I})\bm{y}_{i}}{\sum_{i=1}^{N}\mathcal{N}(\bm{x};s_{t}\bm{y}_{i},s_{t}^{2}\sigma_{t}^{2}\textbf{I})}]$ (19)
Appendix B Additional Experiments
Figure 3: Sample generations from Multistage LDM (CelebA $256\times 256$) and Multistage DPM-Solver (CIFAR-10 $32\times 32$).
B.1 Samples of Additional Generated Results
The generation samples displayed in fig. 3, from the CelebA and CIFAR-10 datasets, demonstrate that our Multistage strategy is capable of producing high-quality results. Notably, it achieves this with reduced training and computational requirements in comparison to baseline methods.
B.2 Comparison with multi-architecture multi-expert (MEME).
In this subsection, we compare our multistage framework with MEME diffusion model [21]. We reimplement the MEME architecture following their released architecture in Table 4 of [21]. We train both our multistage framework and their MEME on the CIFAR-10 dataset and use DPM-Solver for sampling. The results are shown in table 5. From our experiments, Multistage outperforms MEME in FID with much fewer training iterations and total training PFLOPs. However, because [21] only implements the architecture for CelebA-HQ and FFHQ datasets, the method could potentially achieve better results than our reported results with more hyperparameter tunning. Also, the training is unstable and the loss explodes when we try to increase the model capacity of MEME to be comparable to our Multistage.
Table 5: Training and Sampling Efficiency with related works.
Dataset/Method Training Iterations($\downarrow$) GFLOPs($\downarrow$) Total Training PFLOPs($\downarrow$) FID($\downarrow$)
CIFAR-10 $32\times 32$
MEME DPM-Solver [21] $6.0\times 10^{5}$ 13.08 7.85 3.23
Multistage DPM-Solver (Ours) $\textbf{2.5}\times\textbf{10}^{5}$ 18.65 4.66 2.71
B.3 Results in terms of the Precision and Recall Metrics
We evaluate the precision and recall [42] of our proposed architecture on the CIFAR-10 dataset, as results shown in table 6. The recall of the multistage diffusion models is slightly better than their vanilla counterparts. This is potentially because the multistage architecture introduces more degree of freedom when $t=0$ (a separate decoder for this interval), resulting in more diverse data generation.
Table 6: Precision and Recall on CIFAR-10.
METHOD Precision($\uparrow$) Recall($\uparrow$)
DPM-Solver 0.660 0.613
EDM 0.677 0.620
Multistage DPM-Solver 0.657 0.615
Multistage EDM 0.672 0.629
B.4 Ablation Study on the Network Parameters
In this subsection, we perform a series of experiments to determine the best number of parameters within the attempted settings for each stage of our model. For evaluation, we use the best FID scores from all checkpoints. These evaluations are conducted using the DPM-Solver with $20$ Neural Function Evaluations (NFE) on $50,000$ samples. Here, the experimental settings are similar to those used in Section 5.
For the Multistage DPM-solver on CIFAR-10 dataset. The models are trained until they achieve their best FID scores. For calculating FID, we follow the methodology described by [29], using the tensorflow_gan  [^6] https://github.com/tensorflow/gan.
The design of the corresponding parameters for encoders and decoders across different architectures, along with their best FID scores, is detailed in Table 7. We explore various architectures, aiming to maintain a consistent total number of parameters while varying the ratio of shared parameters (encoder parameters) to the total number of parameters. Our ablation study illustrates that Architecture 1 (our chosen model) indeed exhibits the best FID performance.
Architecture Current/Total Parameters GFLOPs ($\downarrow$) FID ($\downarrow$)
1 1/3 169M 29.95 2.35
2/3 108M 17.65
3/3 47M 6.31
Shared 43M 5.75
Total 237M 18.65
2 1/3 168M 27.57 2.9
2/3 138M 21.73
3/3 72M 9.65
Shared 68M 8.98
Total 242M 19.7
3 1/3 189M 29.43 3.1
2/3 160M 23.91
3/3 103M 13.7
Shared 98M 12.93
Total 256M 22.45
4 1/3 174M 26.54 3.34
2/3 126M 17.62
3/3 103M 13.7
Shared 98M 12.93
Total 207M 22.45
Table 7: Ablation study of Multistage DPM-solver on CIFAR-10 dataset. Current/Total: the number of the stage over the total number of stages. Shared: the encoder. Total: combined the encoder and decoders. Total GFLOPS: averaged GFLOPS of all stages weighted by NFE assigned to each stage during sampling of DPM-Solver.
For Latent Diffusion multistage models on CelebA-HQ, models are trained for 500k iterations, and the best checkpoints are used for evaluation. We compute the FID following [8] using the torch-fidelity package [43]. The parameter design and best FID are shown in Table 8. Our ablation study shows the superiority of utilizing Architecture 4 (our chosen model) in terms of the reported FID values.
Architecture Current/Total Parameters GFLOPs ($\downarrow$) FID ($\downarrow$)
1 1/3 238M 80.03 9.37
2/3 189M 64.95
3/3 161M 51.87
Shared 79M 16.69
Total 428M 65.75
2 1/3 206M 72.61 9.73
2/3 158M 57.85
3/3 108M 34.38
Shared 55M 11.6
Total 361M 54.37
3 1/3 274M 88.39 8.43
2/3 224M 72.97
3/3 170M 48.17
Shared 108M 22.71
Total 452M 69.22
4 1/3 316M 105.82 8.38
2/3 224M 72.97
3/3 170M 48.17
Shared 108M 22.71
Total 494M 76.19
Table 8: Ablation study of Multistage LDM on CelebA dataset.
B.5 Ablation Study on the Number of Intervals
In this subsection, we conduct ablation study on the number of intervals of our multistage architecture. We utilize Multistage DPM-Solver training on the CIFAR-10 dataset. To extend algorithm 1, we propose the procedure outlined in algorithm 2 (appendix C). The splitting intervals and Results are given in table 9. We compare the stages from one to five.
Number of Stages Current/Total Interval Parameters GFLOPs ($\downarrow$) FID ($\downarrow$)
1 1/1 [0,1] 108M 17.65 2.75
2 1/2 [0,0.476) 136M 26.59 2.56
2/2 (0.476,1] 95M 16.85
Total 188M 21.49
3 1/3 [0,0.442) 169M 29.95 2.35
2/3 [0.442, 0.630) 108M 17.65
3/3 [0.630,1] 47M 6.31
Total 230M 18.65
4 1/4 [0, 0.376) 159M 29.78 2.67
2/4 [0.376, 0.526) 95M 16.85
3/4 [0.526, 0.726) 71M 11.97
4/4 [0.726, 1] 32M 4.41
Total 284M 17.33
5 1/5 [0, 0.376) 154M 30.38 2.88
2/5 [0.376, 0.476) 88M 16.82
3/5 [0.476, 0.626) 88M 16.82
4/5 [0.626, 0.776) 27M 4.42
5/5 [0.776, 1] 21M 3.28
Total 336M 17.03
Table 9: Ablation study on the number of stages.
B.6 Experimental Setting on Interval Splitting Methods
This section presents the experimental settings of previous timestep clustering algorithms considered in Table 4. We refer readers to the following references [20, 21] for implementation details. Specifically, we compare these methods for the 3 intervals case on CIFAR-10 dataset by training our proposed Multistage architecture with calculated intervals until convergence. The timestep and SNR-based clustering algorithms are relatively easy to implement. For the gradient-based clustering, we collect the gradients of the trained network by the DPM-Solver for clustering. These gradients are generated every 50k training iterations with a total of 450k iterations. We evaluate the performances of these models based on their best FID scores across training iterations.
Appendix C Generalized Alrogithm 1
Algorithm 2 Optimal Denoiser based Timestep Clustering for More Intervals
1:Input: Total samples $K$, optimal denoiser function $\bm{\epsilon}^{*}_{\bm{\theta}}(\bm{x},t)$, interval number $n$, dataset $p_{\text{data}}$
2:Output: Timesteps $t_{1}$, $t_{2}$, $\ldots$, $t_{n-1}$
3:$\mathcal{S}\leftarrow\emptyset$
4:for $k\in\{1,\dots,K\}$ do
5:     $\bm{y}_{k}\sim p_{\text{data}},\bm{\epsilon}_{k}\sim\mathcal{N}(0,\mathbf{I}),t_{a_{k}}\sim[0,1],t_{b_{k}}\sim[0,1]$
6:     $\mathcal{S}^{k}\leftarrow\mathcal{D}(\bm{\epsilon}^{*}_{t_{a_{k}}},\bm{\epsilon}^{*}_{t_{b_{k}}},\bm{y}_{k},\bm{\epsilon}_{k})$
7:     $\mathcal{S}\leftarrow\mathcal{S}\cup\{\left(t_{a_{k}},t_{b_{k}},\mathcal{S}^{k}\right)\}$
8:end for
9:$t_{1},\cdots,t_{i-1}\leftarrow\operatorname*{\text{arg~{}min}}_{t_{1},\cdots,t_{i-1}}\quad\sum\limits_{k=1}^{K}\sum\limits_{i=0}^{n-1}\sum\limits_{t_{i}}^{t_{i+1}}\mathcal{S}^{k}\mathds{1}(t_{a_{k}},t_{b_{k}}\in[t_{i},t_{i+1})),$
[…]
In this section, we design a procedure to extend Algorithm 1 for case of $n\neq 3$. Specifically, we partition the timesteps into $n$ intervals $[0,t_{1}),\ldots,[t_{n-2},t_{n-1}),[t_{n-1},1]$. Intuitively, the intervals need to minimize the summation of the total functional distance within each partition. This corresponds to the following optimization problem.
$\min_{t_{1},\cdots,t_{n+1}}~{}\sum\limits_{k=0}^{n-1}\int\limits_{t_{k}}^{t_{k+1}}\int\limits_{t_{k}}^{t_{k+1}}\mathcal{S}(\bm{\epsilon}^{*}_{t_{a}},\bm{\epsilon}^{*}_{t_{b}})\text{d}t_{a}\text{d}t_{b}\quad\textrm{s.t.}\quad t_{0}=0,~{}t_{n}=1,~{}0<t_{1}<t_{2}<\cdots<t_{n-1}<1\>.$ (20)
In order to solve (20) numerically, we propose Algorithm 2. A solution of the program in step 9 is obtained by discretizeing the time index into $t\in\{0.001,0.026,\dots,0.951,0.976,1\}$. We use the CIFAR10 dataset with $K=50000$ to run the experiment. The following table shows the results of different stage numbers $n$ and different split intervals:
Appendix D Details on Training Procedures
Algorithm 3 Memory efficient training (per GPU)
1:Input: Total Training iterations T, Parameters $\bm{\theta}_{i}$, Interval $[t_{i-1},t_{i}]$ for three stages ($i=1,2,3$), GPU index $k$
2:Output: Parameters $\bm{\theta}^{*}_{i}$
$\text{idx}=k\%3$
4:for $\text{iter}\in\{1,\dots,T\}$ do
5:     Calculate loss and update $\theta_{\text{idx}}$
6:end for
Algorithm 4 Time efficient training (per GPU)
1:Input: Total Training iterations T, Parameters $\bm{\theta}_{i}$, Interval $[t_{i-1},t_{i}]$ for three stages ($i=1,2,3$)
2:Output: Parameters $\bm{\theta}^{*}_{i}$
3:for $\text{iter}\in\{1,\dots,T\}$ do
4:     for $\text{idx}\in\{1,\dots,3\}$ do
5:         Calculate loss and update $\theta_{\text{idx}}$
6:     end for
7:end for
For the training, we employ the same training iterations and loss weights across all stages. Specifically, we propose two algorithms for multi-GPUs parallel training: (i) memory-efficient strategy (algorithm 3) where each GPU updates parameters for a specific stage, requiring less memory, and (ii) time-efficient strategy (algorithm 4) where all GPUs update parameters from all stages. These two algorithms are equivalent from an optimization perspective.